OpenStudy (anonymous):

cot^2 theta/ csc theta = cos theta times cot theta

OpenStudy (anonymous):

whats the answer??? please help!!!!!

OpenStudy (solomonzelman):

\(\Huge\color{green}{ \sf \frac{cot^2θ}{cscθ} =cosθ \times cotθ }\)

OpenStudy (solomonzelman):

we know that \(\Huge\color{red}{ \sf cot^2x =\frac{cos^2x}{sin^2x} ~~~~~~~~~~ AND}\) \(\Huge\color{red}{ \sf cscx =\frac{1}{sinx} }\) OK ?

OpenStudy (anonymous):

ok got it

OpenStudy (solomonzelman):

\(\Huge\color{green}{ \sf \frac{\frac{cos^2θ}{sin^2θ}}{\frac{1}{sinθ}} =cosθ \times cotθ }\) Now, I want to ask you to simplify the fraction on the left. It would be equal to ?

OpenStudy (anonymous):

uhhhhhhh yeah sin ^2 theta???? idk

OpenStudy (anonymous):

i really don't know csc?????

OpenStudy (solomonzelman):

HINT \(\Huge\color{green}{ \sf \frac{\frac{S}{D}}{\frac{A}{B}} =\frac{S}{D} \div \frac{A}{B}=\frac{S}{D} \times \frac{B}{A}}\) \(\Huge\color{green}{ \sf \frac{\frac{cos^2θ}{sin^2θ}}{\frac{1}{sinθ}} =\frac{cos^2θ}{sin^2θ} \div \frac{1}{sinθ}=\frac{cos^2θ}{sin^2θ} \times \frac{sinθ}{1} =?}\)